Circular No. 5810

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET HELIN-LAWRENCE (1993l) E. Helin and K. Lawrence report their discovery of a comet on 0.46-m

Palomar Schmidt telescope films taken by Helin, Lawrence, and M. Nassir. The object appears diffuse with central condensation. The following measurements are by J. Alu:

1993 UT  α_{2000}  δ_{2000}  m_{1}
May 17.43611  16^{h}45^{m}15.^{s}43  −16^{o}05′43′′.1  16.5
17.46510  16 45 14.30  −16 05 44.6
17.47778  16 45 13.75  −16 05 43.6
19.35365  16 44 03.21  −16 07 39.7
19.38368  16 44 02.24  −16 07 44.5

SUPERNOVA 1987A IN THE LARGE MAGELLANIC CLOUD The High Speed Photometer team (R. C. Bless, J. F. Dolan, J. L.

Elliot, E. L. Robinson, G. W. van Citters, J. D. Biggs, P. Boyd, M. J. Nelson, J. W. Percival, M. J. Taylor, and K. Wolinski) report: “The Hubble Space Telescope observed SN 1987A with the High Speed Photometer (+ F160LP filter, 160-700 nm) with a sample time of 100 μs beginning 1992 June 2.45792 UTC for 40 min (73 counts/s), 1992 Nov. 22.80690 for 40 min (89 counts/s), and 1993 Mar. 6.90637 and 6.96637 for 20 min each (100 counts/s). No periodic signal was seen in any data set using an autocorrelation analysis with lags from 400 μs to 0.5 s, or in the power spectrum from the Nyquist frequency (5000 Hz) to 0.01 Hz. Data analysis is still incomplete with respect to techniques and coverage in the frequency domain. Assuming a Crab-like spectrum with similar reddening (cf. Percival et al. 1993, Ap.J. 407, 276), the observed count rates imply a time-averaged magnitude V ∼ 18.2. Comparison with a test signal added to the data indicates that < 1 percent of the flux is pulsed in this spectral bandpass at any of the temporal frequencies examined. This implies an upper limit of mag ∼ 23 in the pulse, using this filter.”

V1974 CYGNI Visual magnitude estimates: Apr. 18.30 UT, 11.4 (W. G. Dillon, Mis-

souri City, TX); 30.99, 11.5 (E. Schweitzer, Strasbourg, France); May 11.91, 11.7 (L. Szentaskó, Budapest, Hungary); 20.95, 12.0 (B. H. Granslo, Fjellhamar, Norway); 29.97, 12.2 (A. Boattini, Florence, Italy).

1993 June 2 Daniel W. E. Green